OpenStudy (sleepyjess):

How can (1/4)x-3=(1/2)x+8 be set up as a system of equations? I got this wrong on the test with an answer of 4y-4x=-12 and 2y-2x=16. I'd like to know what I did wrong and of course, they don't tell you.

OpenStudy (sleepyjess):

\[\frac{ 1 }{ 4 }x-3=\frac{ 1 }{ 2 }x+8\]

OpenStudy (sleepyjess):

@satellite73 ... if you get a chance, can you explain this to me?

OpenStudy (anonymous):

i guess you could write \[\frac{ 1 }{ 4 }x-3=y\\ \frac{ 1 }{ 2 }x+8=y\] but that seem silly

OpenStudy (judygreeneyes):

It does seem silly, but you are probably correct!

OpenStudy (anonymous):

or you could write \[x-4y=12\\ x-2y=-16\] both are equally ridiculous because you have it right in the form to solve

OpenStudy (sleepyjess):

The available answers other than the one I chose were:

OpenStudy (sleepyjess):

4y + 4x = −12 2y + 2x = 16

OpenStudy (anonymous):

i see the problem you multiplied the first one by 4, but \(4\times \frac{1}{4}=x\) not \(4x\)

OpenStudy (sleepyjess):

4y − x = −12 2y − x = 16

OpenStudy (sleepyjess):

4y + x = −12 2y + x = 16

OpenStudy (sleepyjess):

Ok, I see where I got the x part wrong

OpenStudy (anonymous):

\[x-4y=12\\ x-2y=-16\] is what i got multiply both sides by \(-1\) and get \[4y-x=-12\\ 2y-x=16\]

OpenStudy (anonymous):

what a weird arbitrary question

OpenStudy (sleepyjess):

Any idea of why they would do that instead of leaving it as the x-4y and x-2y?

OpenStudy (the_fizicx99):

I've never seen any problem like this. O.e :S

OpenStudy (sleepyjess):

lol @tHe_FiZiCx99

OpenStudy (sleepyjess):

@satellite73 ... I was going to send you a message, but it says you only accept messages from someone you've fanned

OpenStudy (anonymous):

C. 4y + x = −12 2y + x = 16 And D. 4y − 4x = −12 2y − 2x = 16 Are wrong , just took the test